# Bridging Deep Reinforcement Learning and Motion Planning for Model-Free

### v1

@@ IV-C Infeasible Paths Analysis
As shown in Figure 1 (right), compared to the tree-based method, our graph-based method constructs a dense waypoint network that enables comprehensive exploration and facilitates the learning of goal-reaching policies, even in cluttered environments and with unknown dynamics.
V Generalizing Initialization
@@
With the graph-based waypoint covering the whole environment, we could naturally extend the training objective from single start point to arbitrary start points. This brings a desirable property, significant saving computational resource. In an autonomous system with tasks described by temporal logic formulas, if there are $N$ goal regions, in the worst case, if a neural network controller can only move from a specific starting point to a specific endpoint, the system would require $O(N^{2})$ distinct controllers. This indicates that in complex systems, the demand for controllers will increase dramatically as the number of destinations grows. However, as shown in figure 2 under our approach, the required number of controllers is only $O(N)$, i.e., only one controller is needed for each goal region. In this section, we will discuss how our method enables goal-reaching from arbitrary positions within the environment.
Figure 2: Generalize Initialization: We implement our method in the 3D maze environment and present both the top-down view (left) and the result after motion planning (right). The green point denotes the goal, the red point denotes the starting point, and the paths in different colors correspond to the optimal paths originating from different starting points.
V-A Discrete Region Partitioning
In RL training, it is often challenging to guarantee that a trained RL policy capable of completing a task in one initial region can achieve the same performance in another starting region [7, 38]. This limitation typically arises due to the incomplete exploration of the state space by the critic model, leading to underestimation issues. Here, inspired by [39] and [40], we first discretize the geometric space into a set of cell representations (as shown in figure 3).
Figure 3: Discrete Region Partitioning.
Let $\mathcal{E}\subset\mathbb{R}^{d}$ denote a continuous $d$-dimensional environment. We discretize $\mathcal{E}$ by partitioning it into a finite set of non-overlapping cells, each represented as a hypercube of fixed side length $\delta>0$. Formally, the discretized space can be expressed as a grid:
$\mathcal{E}_{\text{dis}}=\{c_{i_{1},i_{2},\dots,i_{d}}\mid i_{k}\in\mathbb{Z},\forall k=1,\dots,d\}$ (4)
where each cell $c_{i_{1},i_{2},\dots,i_{d}}$ is defined as
$c_{i_{1},i_{2},\dots,i_{d}}=\left\{x\in\mathbb{R}^{d}\mid i_{k}\delta\leq x_{k}<(i_{k}+1)\delta,\forall k=1,\dots,d\right\}.$ (5)
Here, $\delta$ is referred to as the spatial granularity, which determines the resolution of the discretization. The choice of $\delta$ has a certain impact on our method. However, for simplicity, we set $\delta$ as one-tenth of the smallest dimension length of the environment, providing a reasonable discretization resolution. Also we slightly abuse the notation $c_{i}$ as $c_{i_{1},i_{2},\dots,i_{d}}$, where $i$ is a shorthand for the tuple $(i_{1},i_{2},...,i_{d})$.
One intuitive advantage of partitioning the environment into discrete cells is that it significantly reduce the number of waypoints we need to iterate in the training process. Even in low-dimensional space with the same waypoint density, graph-based method generates significantly more waypoints and connections than tree-based methods. And this situation would be more extreme while we scale up the environment. By assigning each waypoint to its corresponding cell $c_{i}$, we can restrict the search within a much smaller region, significantly reducing the search space while maintaining efficient path planning. Therefore, here we extend the $G=(V,E,Cost)$ to $G=(V,E,Cost,Cell)$, where $Cell:\mathcal{S}\rightarrow\mathcal{E}_{dis}$ is a function mapping from geometry space to the discrete cell index.
V-B Weighted Sampled Starter
A common application of our approach is providing controllers for autonomous systems. Our method has already established reward guidance across every region of the environment. However, relying solely on randomly sampled starting points often fails to effectively explore all areas of the map and does not guarantee reachability from those regions to the goal. This limitation significantly impacts practical applications. For instance, in [41], verification results show that the set of controllers may fail to accomplish the specified task.
Recall the discretized state space be defined as $\mathcal{E}_{\text{dis}}=\{c_{i}\mid i\in\mathcal{I}\}$, where each $c_{i}$ represents a fixed-size grid cell. We define $N_{i}$ as the number of times the agent visits cell $c_{i}$ and compute the exploration coverage as $P_{i}=\frac{N_{i}}{\sum_{j\in\mathcal{I}}N_{j}+\epsilon}$ where $\epsilon$ is a small positive constant to prevent division by zero. To prioritize under-explored regions, we define the sampling weight: $w_{i}=\frac{1}{P_{i}+\epsilon}$, which is then normalized as $\tilde{w}_{i}=\frac{w_{i}}{\sum_{j\in\mathcal{I}}w_{j}}$. Using the normalized weights $\tilde{w}_{i}$, we sample a starting grid cell $c_{\text{start}}$ from a categorical distribution $c_{\text{start}}\sim\text{Categorical}(\mathcal{I},\tilde{w})$. Within the selected grid cell $c_{\text{start}}$, the exact starting state is uniformly sampled: $s_{\text{start}}\sim\mathcal{U}(c_{\text{start}})$.
As training progresses, visited states contribute to updating the visit counts of their respective grid cells, thereby gradually reducing the weight assigned to previously explored areas. Specifically, at the end of each training episode. For the trajectory $\tau=\{s_{0},\dots,s_{T},\}$ in the same episode, we first mapping from geometry space to cell space $\tau_{c}={Cell(s_{t}),\ \forall t\in[0,T]}$. Then we construct an unduplicated sets $\mathcal{C}_{visited}=\bigcup^{T}_{t=0}\{Cell(s_{t})\}$. The visit count for each explored cell $c_{i}$ is updated as: $N_{i}\leftarrow N_{i}+1,\quad\forall i\text{ where }c_{i}\in\mathcal{C}_{visited}$ This approach dynamically adjusts the start state distribution, ensuring thorough exploration and improving the coverage of the learned policy across diverse initializations.
VI Experimental Results
@@
We validated our framework with different dynamic models. The experiments highlight that our method demonstrates superior performance in environments with disturbances where the environment changes after motion planning. We evaluated both our method and the baseline approaches using widely adopted reinforcement learning algorithms, i.e., PPO. We use high quality implementation from [42], mostly with default hyperparameters without further tuning. The results show that our approach is able to avoid potentially high-risk regions and significantly improves the success rate.
Figure 4: 2d and 3d navigation in environments with unmodeled disturbance, where light blue area, brown area and gray area represents wetted area, muddy area and strong wind area.
Figure 5: Comparison with other reward functions and validation of weighted sample start in 2D dynamics.
Figure 6: Comparison with other reward and the validation of weighted Sample Start in 3d dynamics
Baseline Approaches: The experiments were setup in both 2d and 3d cluttered environments where we compared our methods with others in different scales. We compare our method against the RRT-guided reward [24] and the binary reward to evaluate its ability to navigate cluttered environments with disturbances. Additionally, we examine the impact of the weighted sampled start strategy by comparing it with the standard approach without weighted sampling.
Quadrotor Model: We first implement the car-like model of Pybullet[43] physical engine shown in Figure 5. In this environment, we designed the obstacle layout to be highly cluttered, resembling a maze, while leaving a relatively simple path available-i.e., bypassing the maze from the upper corner.
We then blocked this simple path to demonstrate two key aspects of our method: it not only ensures successful goal-reaching but also naturally selects routes with lower potential risk. Compared to tree-based methods, which are sensitive to disturbances and unable to adjust their trajectory based on risk-being constrained to follow the pre-planned path-our approach exhibits significantly greater robustness. In this experiments, when we increased the disturbance level to the point that the RRT path was directly broken and the method failed entirely, our approach still performed reliably and successfully reached the goal. Meanwhile, we observed that the weighted sample start significantly improves the convergence speed compared to random initialization.
Autonomous Vehicle: We then implement the car-like model of Pybullet[43] physical engine shown in Figure 6. We randomly generated obstacles and introduced regions with strong winds, mud, and water along the path, representing lateral thrust, significant speed reduction, and impassable terrain, respectively. Similarly to the quadrotor scenario, when both binary reward and tree-based reward methods failed to complete the task, our framework still performed reliably. However, we did not observe a significant advantage of the weighted sample start over random start.
@@
Success Certificate: We statistically run 200 trials of the learned policies for different selected start position and record the average success rates of both models, i.e., autonomous vehicle and quadrotor. The results are shown in Table I. We observe that in cluttered environments, assuming that all of the planned path is feasible and well tracked, the success rates of the binary reward is 0, while our method and tree-based reward achieve 100%; In cluttered environments which contain infeasible region, the success rates of all other baselines are 0, and our method achieves success rates near 100%. As a result, the robustness of the performance has improved significantly in response to environmental challenges.
TABLE I: Analysis of success rates.
Path Dynamic model Baseline rate Success rate
Feasible Quadrotor RRG 100%
Only RRT 100%
Binary 0%
Feasible Quadrotor RRG 100%
Only RRT 100%
Binary 0%
Contains Vehicle RRG 100%
Infeasible RRT 0%
Region Binary 0%
Contains Vehicle RRG 100%
Infeasible RRT 0%
Region Binary 0%
VII Conclusion
@@
DRL suffers from inefficient exploration, especially in cluttered environments. Compared to random exploration, motion-planning-based methods can provide effective guidance for learning. However, tree-based planners that yield a single path may fail to adapt when environmental changes, limiting their guidance. This paper proposes a novel graph-based reward scheme that maintains a model-free setup while offering global guidance coverage to encourage optimal path exploration, and significantly reduce the number of controllers required. Our approach fully leverages the advantages of model-free RL and guarantees the goal-reaching performance for the original task. Future works include Sim2Real generalization and the extension to multi-agent collaboration.

### v2

@@ 4.3 Infeasible Paths Analysis
As illustrated in Figure 2 (right), in a complex but highly connected maze, our graph-based method can easily identify alternative paths to the goal. Even from nearby but distinct start positions, the agent successfully reaches the same goal via different trajectories, demonstrating the diversity and flexibility of path selection enabled by dense connectivity.
4.4 Advantages and Extensions
With the graph-based waypoint covering the whole environment, we could naturally extend the training objective from single start point to arbitrary start points. This brings some desirable properties that enable multiple extensions.
@@ 4.4 Advantages and Extensions
Scalable Control: In an autonomous system with tasks described by temporal logic formulas, if there are $N$ goal regions, in the worst case, if a neural network controller can only move from a specific starting point to a specific endpoint, the system would require $O(N^{2})$ distinct controllers. This indicates that in complex systems, the demand for controllers will increase dramatically as the number of destinations grows. A similar limitation is discussed in [24], where a separate controller is trained for each unique start-goal pair. However, by generalizing the initialization, the required number of controllers is only $O(N)$, i.e., only one controller is needed for each goal region. This generalization is particularly beneficial for LTL-specified tasks, a representative setting in cluttered environments, as it substantially mitigates onboard resource limitations and enhances system scalability.
Handling Unseen Situations: The ability to reach goals from arbitrary initializations can be further strengthened by integrating recovery mechanisms, such as safety filters—preferably model-free ones—as well as fallback policies or local replanning modules, to handle unexpected disturbances and transform otherwise infeasible scenarios into solvable tasks. This capability provides a significant advantage over prior methods, inherently enhancing system robustness and supporting generalization to new scenarios. Although recovery is not the main focus of this work, we include a simple distance-based filter in our experiments to illustrate the concept.
5 Experimental Results
@@
We evaluated both our method and the baseline approaches using widely adopted reinforcement learning algorithms, i.e., PPO. We use a high-quality implementation of Raffin et al. [30], with most hyperparameters left at their default settings. The results show that our approach significantly improves the success rate and training efficiency.
Figure 3: Comparison of training efficiency (consumption) across different methods. Each plot shows the mean and standard error over 10 runs. Methods that do not appear in certain parts of the curves indicate overlap at the bottom.
Baseline Approaches: We conducted experiments in both 2D and 3D cluttered environments across different scales. To benchmark our approach, we compared against representative methods from count-based and curiosity-driven exploration, namely ICM [9] and RND [10] in cluttered environments. Additionally, we evaluated our method against the RRT-guided reward [24] and a distance-based reward (referred to as binary) to assess its robustness in navigating under disturbances.
Dynamic Model: We evaluate our method across two representative dynamic models implemented using the PyBullet physics engine [31], corresponding to distinct control modes. The first is a quadrotor model controlled via velocity commands in 3D space (i.e., $[v_{x},v_{y},v_{z}]$), while the second includes a car-like ground vehicle and a quadruped robot, both controlled using high-level commands in planar space (i.e., $[v_{x},\dot{\psi}]$). While accurate modeling and low-level control of these systems are not the focus of this work, such components can be readily obtained from prior work on robust locomotion and dynamics learning [32, 33], or replaced with task-specific pretrained controllers depending on the target platform.
Figure 4: Comparison with other reward functions. The bar height (lower is better) indicates the number of steps required for the agent to complete the task for the first time during training.
Task Design: We consider a goal-reaching task where the agent must navigate from various initial positions to a designated goal region. Both the 2D and 3D environments are structured as mazes, where vanilla distance-based rewards are insufficient for effective navigation. Further, the 3D environment is constructed as a $4\times 4\times 4$ room grid, where each floor contains only a single passage leading to the next level, and rooms within the same floor are interconnected by doors. The goal is placed at the topmost layer, while starting positions are uniformly sampled across the lower three floors. Additionally, for comparisons with the RRT-based method [24], we randomly block three doors on the same floor to simulate scenarios where path planning becomes infeasible.
Efficient Exploration: We consider each training task as a compositional reinforcement learning problem, where the agent must learn reusable primitives that enable reaching a goal from any designated region within the environment. Training is deemed successful once the agent reaches the goal region from all ten predefined test starting positions. To evaluate learning efficiency, we define the consumption as the global step at which this success criterion is first achieved and the total training budget. A lower consumption indicates more efficient exploration. If the agent fails to meet the success condition within the allotted training steps, a consumption value of highest global steps is assigned. As shown in Fig. 3 and Fig. 4, although the agent without guidance is occasionally able to reach the goal from certain starting positions, nearly all of them fail to satisfy the overall task criterion, which is reflected in their exceedingly high step consumption.
@@
Success Certificate: We run 200 trials of the learned policies for different selected start positions and evaluate the performance with both dynamic models. The results are shown in Table 1(b). We observe that in cluttered environments, assuming that all of the planned path is feasible and well tracked, the success rates of all the other rewards are 0, while our method and tree-based reward achieve 100%; In cluttered environments that contain infeasible regions, all the baseline methods fail with success rates dropping to 0, while our approach consistently achieves near 100% success. Notably, for the methods relying on RRT-generated reward, once the planned path is blocked by obstacles, the reward structure effectively degrades into a sparse, binary signal, making learning extremely difficult in such settings.
Table 1: Success analysis under different path feasibility settings.
(a) Feasible-only paths
Method Quadrotor Vehicle
RRG
RRT
Others
(b) Paths containing infeasible regions
Method Quadrotor Vehicle
RRG
RRT
Others
Generalization to Unseen Situations: As discussed in Sec. 4.4, a key advantage of our framework lies in its ability to generalize to arbitrary start states and integrate recovery mechanisms seamlessly. To evaluate this capability, we conduct a visualization experiment as shown in Fig. 5. The left panel displays the discounted success score, defined as $\mathbb{I}_{\text{success}}\cdot(\text{steps})^{\gamma}$, where $\mathbb{I}_{\text{success}}$ is the indicator function for goal-reaching and $\gamma\in(0,1)$ is a discount factor that emphasizes earlier success. This illustrates that the agent learns a globally coherent value landscape that supports goal-reaching from diverse initializations. In the middle panel, we visualize 10 representative trajectories from randomly selected start points, all successfully reaching the goal without any external intervention. To simulate realistic failures or unexpected conditions, we introduce unseen obstacles in the test environment and activate a simple distance-based safety filter to avoid entering unsafe zones. The resulting trajectories, shown in the right panel, demonstrate that the agent can dynamically adapt its path and still reach the goal, confirming the effectiveness of our method in handling perturbed and previously unseen situations.
Figure 5: Demonstration of generalization and recovery. Left: Q-map showing learned goal proximity across the state space. Middle: Trajectories from 10 random start positions reaching the goal in an undisturbed environment. Right: Recovery behavior with unseen obstacles and a distance-based safety filter.
6 Conclusion
@@
DRL suffers from inefficient exploration, especially in cluttered environments. Compared to random exploration, motion-planning-based methods can provide effective guidance for learning. However, tree-based planners that yield a single path may fail to adapt when environmental changes, limiting their guidance. This paper proposes a novel graph-based reward scheme that maintains a model-free setup while offering global guidance coverage to encourage optimal path exploration, and significantly reduce the number of controllers required. Our approach fully leverages the advantages of model-free RL and guarantees the goal-reaching performance for the original task.
Appendix
Appendix A Cost-to-Go Calculation for Planning-Based Rewards
Before discussing the algorithms in detail, we introduce the following algorithmic primitives used in algorithms below:
Discretization
The function DiscretizeEnv divides the continuous environment $\mathcal{X}$ into a uniform grid. Each cell is classified as either free or occupied, depending on whether it overlaps with an obstacle. The resulting map structure serves as the input graph for A* search.
Neighbors
Given a state $s$ and a map structure Map, the function Neighbors$(s,\text{Map})$ returns the set of neighboring states that are directly reachable from $s$. In a 4-connected grid, these are the adjacent horizontal and vertical cells; in an 8-connected grid, diagonal neighbors are also included. In a graph like RRG, neighbors are determined by edges in $E$.
Collision check
The function isObstacle$(s)$ returns true if the state $s$ is within an obstacle region, and false otherwise. This function is used to prune invalid neighbors from expansion.
Cost
The function Cost$(s,s^{\prime})$ returns the traversal cost between two adjacent states $s$ and $s^{\prime}$. This is typically set to the Euclidean or Manhattan distance between the states for grid maps, or to edge weights in roadmap graphs.
Heuristic
The function $h(s)$ provides an admissible estimate of the cost from state $s$ to the goal state $s_{\text{goal}}$. Common choices include Manhattan or Euclidean distances in grid settings.
Path reconstruction
Once the goal state is reached, the function ReconstructPath$(s_{\text{goal}})$ traces backward through the parent pointers to construct the shortest path from $s_{\text{start}}$ to $s_{\text{goal}}$. The result is a trajectory $\mathbf{X}^{*}=\{s_{0},s_{1},\dots,s_{d}\}$ such that $s_{0}=s_{\text{start}}$ and $s_{d}=s_{\text{goal}}$.
Algorithm 1 RRG with Dijkstra Distance Calculation
1:Initialize $G=(V,E)$ with start node $s_{\text{start}}$
2:for $i=1$ to $N$ do
3:     $s_{\text{rand}}\leftarrow\text{Sample()}$()
4:     $s_{\text{nearest}}\leftarrow\text{Nearest}(V,s_{\text{rand}})$
5:     $s_{\text{new}}\leftarrow\text{Steer}(s_{\text{nearest}},s_{\text{rand}})$
6:     if $\text{CollisionFree}(s_{\text{nearest}},s_{\text{new}})$ then
7:         $V\leftarrow V\cup\{s_{\text{new}}\}$
8:         $E\leftarrow E\cup\{(s_{\text{nearest}},s_{\text{new}})\}$
9:         $S_{\text{near}}\leftarrow\text{Near}(V,s_{\text{new}})$
10:         for all $s_{\text{near}}\in S_{\text{near}}$ do
11:              if $\text{CollisionFree}(s_{\text{near}},s_{\text{new}})$ then
12:                  $E\leftarrow E\cup\{(s_{\text{near}},s_{\text{new}})\}$
13:              end if
14:              if $\text{CollisionFree}(s_{\text{new}},s_{\text{near}})$ then
15:                  $E\leftarrow E\cup\{(s_{\text{new}},s_{\text{near}})\}$
16:              end if
17:         end for
18:     end if
19:end for
20:for all $s\in V$ do
21:     $d(s)\leftarrow\text{Dijkstra}(G,s,s_{\text{goal}})$triangleright obtain the distance to the goal
22:end for
As shown in Algorithm 1, the RRG algorithm begins by initializing the graph $G=(V,E)$ with the start node $s_{0}$ and no edges (line 1). In each iteration, a sample $s_{\text{rand}}$ is drawn, and its nearest neighbor $s_{\text{nearest}}$ is found (lines 3–4). A new state $s_{\text{new}}$ is generated by steering toward the sample (line 5), and if the path is valid, it is added to the graph along with edges to nearby nodes (lines 6–14). Once the graph is constructed, Dijkstra’s algorithm (as shown in Algorithm 2) is executed to compute the cost-to-go values from each node to the goal $s_{\text{goal}}$ (lines 20–21).
Algorithm 2 Dijkstra’s Algorithm for Cost-to-Go Computation
1:function Dijkstra($G,s_{\text{goal}}$)
2:     for all $s\in V$ do
3:         $d(s)\leftarrow\infty$
4:     end for
5:     $d(s_{\text{goal}})\leftarrow 0$
6:     Initialize priority queue $Q\leftarrow\{s_{\text{goal}}\}$
7:     while $Q$ is not empty do
8:         $s_{\text{current}}\leftarrow\text{ExtractMin}(Q)$
9:         for all $s_{\text{neighbor}}$ in Neighbors($s_{\text{current}}$do
10:              $c\leftarrow\text{Cost}(s_{\text{current}},s_{\text{neighbor}})$
11:              if $d(s_{\text{neighbor}})>d(s_{\text{current}})+c$ then
12:                  $d(s_{\text{neighbor}})\leftarrow d(s_{\text{current}})+c$
13:                  Update $Q$ with $d(s_{\text{neighbor}})$
14:              end if
15:         end for
16:     end while
17:     return $d(\cdot)$
18:end function
In addition to sampling-based planning methods such as PRM and RRG described in the main text, we also consider a sample-free alternative using a discrete grid map. In this setting, the environment is discretized into a uniform grid, and A* search is directly applied to compute feasible waypoint sequences. This approach avoids the use of sampling and graph construction, and instead leverages the regular connectivity of the grid. The complete procedure is shown in Algorithm 3
Using the map either generated by sampling based method like RRG, or use the discrete map mentioned above, Algorithm 3 performs a standard A* search to compute the shortest path from a start state $s_{\text{start}}$ to a goal state $s_{\text{goal}}$. The algorithm initializes the open set with $s_{\text{start}}$ and assigns its cost-to-come $g(s_{\text{start}})=0$ and heuristic value $f(s_{\text{start}})$ (lines 2–4). At each iteration, the node with the smallest $f$ value is selected for expansion (line 5). If the goal is reached, the optimal path is reconstructed and returned (lines 7–8). Otherwise, the algorithm expands the current node $s_{\text{current}}$ by iterating over its neighbors as defined by the map (line 10). Invalid neighbors, such as those in obstacles or already explored, are skipped (line 13). For valid neighbors, the tentative cost is computed, and if a better path is found, the corresponding $g$, $f$, and parent values are updated (lines 16–22). This process continues until the goal is found or the open set is exhausted. If the goal is unreachable, the algorithm returns failure (line 25).
Algorithm 3 A* Search on General Map
1:function A*($\text{Map},s_{\text{start}},s_{\text{goal}}$)
2:     OpenSet $\leftarrow\{s_{\text{start}}\}$
3:     $g(s_{\text{start}})\leftarrow 0$
4:     $f(s_{\text{start}})\leftarrow g(s_{\text{start}})+h(s_{\text{start}})$
5:     while OpenSet is not empty do
6:         $s_{\text{current}}\leftarrow\text{argmin}_{s\in\text{OpenSet}}f(s)$
7:         if $s_{\text{current}}=s_{\text{goal}}$ then
8:              return ReconstructPath$(s_{\text{goal}})$
9:         end if
10:         Remove $s_{\text{current}}$ from OpenSet
11:         Add $s_{\text{current}}$ to ClosedSet
[…]
13:              if $s_{\text{neighbor}}\in$ ClosedSet or isObstacle then
14:                  continue
15:              end if
16:              $g_{\text{tent}}\leftarrow g(s_{\text{current}})+\text{Cost}(s_{\text{current}},s_{\text{neighbor}})$
17:              if $s_{\text{neighbor}}$ not in OpenSet or $g_{\text{tent}}<g(s_{\text{neighbor}})$ then
18:                  $g(s_{\text{neighbor}})\leftarrow g_{\text{tent}}$
19:                  $f(s_{\text{neighbor}})\leftarrow g(s_{\text{neighbor}})+h(s_{\text{neighbor}})$
20:                  Parent$(s_{\text{neighbor}})\leftarrow s_{\text{current}}$
21:                  Add $s_{\text{neighbor}}$ to OpenSet
22:              end if
23:         end for
24:     end while
25:     return $\infty$ $\triangleright$ failure
26:end function
Appendix B Theorem 1 proof
Proof.
We first prove this reward is non-Markovian in the original state space $\mathcal{S}$. With the definition of MDP [2], we prove it by a contradiction. Assume, for the sake of contradiction, that the reward function in the original state space is Markovian. That is, at each timestep $t$
$r(s_{t},a_{t},h_{t-1})=r(s_{t},a_{t})$ (3)
It indicates the Markovian reward is only related to the current state and action. Suppose the agent has already achieved one r-ball, thereby updating its internal record of $D_{min}$ (which keeps track of exploration progress). Then the agent moves back and achieves this r-ball again. For these two visits, the reward is obviously different since in the second time, the agent visits the same r-ball with the same $D_{min}$, which would not lead to a reward. In other words, there exist two different histories $h_{1}$ and $h_{2}$ leading to the same state-action pair $(s,a)$ but resulting in different values of previous $D_{min}$, causing $r(s,a,h_{1})\neq r(s,a,h_{2})$. This contradicts the Markov property, proving that the reward in the original state space is inherently non-Markovian.
To address the above contradiction, we incorporate $D_{min}$ into the agent’s observable state. Specifically, we extend the original state space $\mathcal{S}$ to
$\mathcal{S}^{\times}\;=\;\bigl{\{}(\,s,\,D_{min}\!)\;\mid\;s\in\mathcal{S},\ D_{min}\in\mathbb{R}_{\geq 0}\bigr{\}}.$
In other words, each state in $\mathcal{S}^{\times}$ explicitly tracks both the physical state $s$ and the most up-to-date information $D_{min}$ needed for the reward function. In this augmented space, all historical information necessary for computing the motion-planning-guided reward is encapsulated in the tuple $(s,D_{min})$. As a result, whenever the agent revisits a physical location, the question of whether a specific $r$-ball has been “newly achieved” or“already achieved” is resolved by the corresponding $D_{min}$ values stored in the augmented state. Consequently, the reward function $\tilde{r}$ can be written as: $\tilde{r}\bigl{(}(s_{t},D_{min}),\,a_{t}\bigr{)}$, or, if needed, $\tilde{r}\bigl{(}(s_{t},D_{min}),\,a_{t},\,(s_{t+1},D_{min}^{\prime})\bigr{)}$, thereby depending only on the current augmented state (and possibly the next state), making it Markovian.
Formally, for any pair of histories $h_{1}$ and $h_{2}$ that arrive at the same augmented state $(s_{t}^{\times})$ and take the same action $a_{t}$, the updated $(s_{t+1}^{\times})$ remains the same (in distribution) and so does the associated reward. This is precisely the Markov property in the augmented space:
No additional information from the entire trajectory $h_{t-1}$ is necessary once $D_{min}$ is encoded in the state. ∎
Appendix C Proposition 1 Proof
Proof.
$\hat{r}_{t}(s_{t})=\left\{\begin{array}[]{cc}R_{-},&\text{if $s_{t}$ is in theobstacles},\\ R_{++},&\text{if $s_{t}$ is in the goal area }\\ 0,&\text{otherwise,}\end{array}\right.$ (4)
Let the original reward $\hat{r}$ be defined as in Eq. (4), and the augmented reward $\tilde{r}$ be defined as in Eq. (2).
We first show that there exists an optimal policy that reaches the goal under the original reward $\hat{r}$. For the MDP we described in Section 3, the existence of such a policy is guaranteed by standard results in measure-theoretic Markov decision processes (see, e.g., Theorem 3.2 in [34] and Theorem 8.7.2 in [26]).
Then we adopted Theorem 1 in [35], which shows that the optimal policy remains invariant if the reward transformation is of the form:
$R^{\prime}(s,a,s^{\prime})=R(s,a,s^{\prime})+\gamma\Phi(s^{\prime})-\Phi(s)$
where $\Phi:\mathcal{S}\rightarrow\mathbb{R}$ is a potential function defined over the state space $\mathcal{S}$, and $\gamma$ is the discount factor. To satisfy the theorem, it suffices to construct such a potential function $\Phi(s)$ that renders the augmented reward a potential-based transformation.
This condition does not hold in our case if the potential function depends on historical information such as the previous minimum distance to the goal. However, the augmented state space includes the historical minimum distance $D_{min}$, ensuring that the potential function $\Phi(s^{\times})$ remains a well-defined function over states. Here $D_{min}(\tau_{t})=D_{min}^{t}$.
In our framework, we could explicitly construct a potential function:
$\Phi(s_{t}^{\times})=\begin{cases}C(1-e^{-k(d_{0}-D_{min}^{t})}),&\text{if }D_{min}^{t}\leq d_{0}\\ 0,&\text{otherwise}\end{cases}$
where $d_{0}=D_{min}^{t-1}$ is the minimum distance to the goal achieved so far (included in the state representation), $C$ is a constant, and $k$ is a sharpness parameter.
Let $C=R_{+}$, $\gamma=1$ and choose large $k\gg 0$. Define the shaping term as $F(s^{\times},a,s^{\times^{\prime}})=\gamma\Phi(s^{\times^{\prime}})-\Phi(s^{\times})$. In the case $d^{\prime}<d$, which means the agent is closer to the goal, we get $F(s^{\times},a,s^{\times^{\prime}})\approx R_{+}$; if $d^{\prime}\geq d$, which means the agent stays or gets further from the goal, then $F(s^{\times},a,s^{\times^{\prime}})\approx 0$. Then we have the reward we defined $\tilde{r}(s_{t}^{\times})=\hat{r}(s_{t}^{\times})+F(s^{\times},a,s^{\times^{\prime}})$, which is exactly in the form required by the theorem 1 of [35]. Hence, it does not alter the optimal policy of the original MDP.
Eventually, we show that the augmented reward accelerates convergence. Temporal-Difference (TD) learning is a core component of almost all modern reinforcement learning algorithms, serving as the basis for value updates during training. The TD error at time step $t$ is defined as:
$\delta_{t}=r(s_{t}^{\times})+\gamma V(s_{t+1}^{\times})-V(s_{t}^{\times}),$
which serves as the core learning signal for value updates. Let the augmented reward be $\tilde{r}=\hat{r}+\gamma\Phi(s^{\times^{\prime}})-\Phi(s^{\times})$. The TD error under shaping becomes:
$\delta^{\prime}_{t}=\hat{r}(s_{t}^{\times})+\gamma\Phi(s_{t+1}^{\times})-\Phi(s_{t}^{\times})+\gamma V(s_{t+1}^{\times})-V(s_{t}^{\times}).$
In sparse-reward settings, most transitions yield $\hat{r}=0$, and thus the shaping term dominates the learning signal. If $\Phi$ is positively correlated with goal proximity, then $\delta^{\prime}_{t}$ reflects meaningful progress even when the environment provides no extrinsic reward. According to the convergence theory of TD learning [2], larger TD errors, which remain proper learning rates, lead to faster value function updates. Therefore, our augmented rewards enhance early-stage TD signals and accelerate convergence in sparse environments.